OpenStudy (anonymous):

I'm trying to solve this but I honestly have no idea how to move forward. All I have done so far is this: theta=tan^-1(-1) Now I'm stuck <_>

OpenStudy (ash2326):

@greysmither what's tan 45?

OpenStudy (anonymous):

it is 135 degree ,, Aren't u allowed to use calculator??

OpenStudy (ash2326):

Avva 135 is just one of the answer, it has several answers

OpenStudy (amistre64):

tan-1 only has one answer :)

OpenStudy (ash2326):

@amistre64 tan (theta)=-1 will have several solutions

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

yup you can add 180 degrees and you'll figure out that 135 n 315 both have tan of -1

OpenStudy (ash2326):

\[\tan \theta=-1\] so \[\theta = n\pi +3\frac{\pi}{4}\] n=0, 1, 2, 3 and so on

OpenStudy (anonymous):

So this pretty much a unit circle-dependent problem?

OpenStudy (anonymous):

this is*

OpenStudy (ash2326):

Yeah:)

OpenStudy (anonymous):

Could it also be \[7\pi/4 + \pi n\]?

OpenStudy (anonymous):

since both of those quadrants create -1?

OpenStudy (ash2326):

Yeah it should be \[ n\pi \pm \pi/4\] n=1,2 ,3 and so on I missed that, it'd include \(7\pi/4 + n\pi\) also

OpenStudy (anonymous):

I thought that you couldn't include pi/4 and 5pi/4 in quadrants 1 and 3 because they equal 1 when I'm looking for -1?

OpenStudy (anonymous):

So confusing :(

OpenStudy (ash2326):

Sorry, Just wait for some time, I'll help you:)

OpenStudy (ash2326):

Sorry kept you waiting:(

OpenStudy (anonymous):

Its okay I really appreciate your help

OpenStudy (ash2326):

We have \[\tan x=-1\] so x is eithere 3pi/4 , 7pi/4 and so on \[x= n\pi -\pi/4\] if n=1 x=3 pi/4 n=2 x= 7 pi/4 so this will include all possible solutions. This is your answer

OpenStudy (anonymous):

So, just to clarify \[x=-\pi/4+n \pi\] is the final answer. I can't thank you enough!! I wish I could give you 100 medals ;)

OpenStudy (ash2326):

Thanks Greysmither , this is the answer:)